### v2

# Lower Bounds for Faithful Linear Representations of Subgroups of the Mapping Class Group

Abstract.
Recently, Korkmaz established the lower bound of 3g23g-2 for the dimension of a faithful representation of the mapping class group of an orientable surface of genus $g\geqslant 3$. We raise this bound to 4g34g-3 in the setting of surfaces of genus $g\geqslant 7$. A new ingredient is a finer study of the commutation relations in $\operatorname{PMod}(\Sigma)$. We use the relations arising from a certain pants decomposition of $\Sigma_{g}$ to show that any representation of dimension $\leqslant 4g-4$ is forced to kill a natural subgroup of the Torelli group.
We also establish lower bounds for the dimension of faithful representations of related groups: the Johnson group of a closed surface, arbitrarily low terms of the Johnson filtration of a compact surface with one boundary component and pure braid groups. These lower bounds grow linearly on the genus of the surfaces and the number of strands of the braids. Finally, we also provide some evidence that greater lower bounds for the low-genus cases should lead to improved lower bounds for $g\gg 0$.
2020 Mathematics Subject Classification:
Primary 57K20; Secondary 20F38 15A30
1. Introduction
Denote by $\Sigma=\Sigma_{g,r}^{b}$ the compact connected orientable surface of genus $g$ with $b$ boundary components and $r$ marked points $P=\{x_{1},\ldots,x_{r}\}$ in its interior. Its mapping class group is the group $\operatorname{Mod}(\Sigma)=\pi_{0}\operatorname{Diff}^{+}(\Sigma,\partial\Sigma)_{P}$ of orientation-preserving self-diffeomorphisms of $\Sigma$ fixing the boundary point-wise and permuting the marked points, up to isotopies. The pure mapping class group $\operatorname{PMod}(\Sigma)\leqslant\operatorname{Mod}(\Sigma)$ is the subgroup of mapping classes fixing $P$ point-wise. Let $\Sigma_{g}=\Sigma_{g,0}^{0}$.
The theory of mapping class groups plays a central role in low-dimensional topology, as closed $3$-manifolds may be encoded by elements of $\operatorname{Mod}(\Sigma_{g})$ via mapping tori or Heegaard splittings. The group $\operatorname{PMod}(\Sigma)$ is also of importance in algebraic geometry, as it may be seen as the (orbifold) fundamental group of the moduli space of complex curves.
Yet, basic questions about its linear representations $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ remain unanswered. Recent years have seen intense activity around the study of low-dimensional representations of $\operatorname{PMod}(\Sigma)$, as well as the question of linearity of such groups – which is currently open for $g\geqslant 3$.
Improving results of Funar [13] and Franks–Handel [12], Korkmaz [23] showed any $d$-dimensional linear representations of $\operatorname{PMod}(\Sigma)$ with $d<2g$ is trivial for $g\geqslant 3$. Korkmaz then went on to show that, for $g\geqslant 3$, any nontrivial $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{2g}(\mathbb{C})$ is conjugate to the so-called symplectic representation $\Psi:\operatorname{PMod}(\Sigma)\to\operatorname{Sp}_{2g}(\mathbb{Z})$. When $\Sigma$ has no marked points, its kernel is a normal subgroup of primary importance, known as the Torelli subgroup $\mathcal{I}(\Sigma)\leqslant\operatorname{PMod}(\Sigma)$.
Denoting by $d(G)$ the smallest dimension of a faithful linear representation of a group $G$ and setting $d(\Sigma)=d(\operatorname{PMod}(\Sigma))$, Korkmaz also established $d(\Sigma)\geqslant 3g-2$ for $g\geqslant 3$. He showed that, when $g\geqslant 3$ and $m\leqslant g-3$, any $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{2g+m}(\mathbb{C})$ is forced to kill the $m^{\text{th}}$ derived subgroup $K_{\Sigma^{\prime}}^{(m)}$ of a certain subgroup $K_{\Sigma^{\prime}}\leqslant\mathcal{I}(\Sigma)$, where $\Sigma^{\prime}\subseteq\Sigma$ is a genus $3$ subsurface.
Kasahara [18] classified all $(2g+1)$-dimensional representations for $g\geqslant 7$. Recently, Kaufmann–Salter–Zhang–Zhong [19] further improved Korkmaz’ results by classifying all $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d\leqslant 3g-3$ in the $g\geqslant 4$ and $b+r\leqslant 1$ setting, showing that any such representation is conjugate to the direct sum of a $(2g+1)$-dimensional representation with copies of the trivial representation $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{1}(\mathbb{C})$.
Their result shed light into Korkmaz’ lower bound of 3g23g-2 in the setting of closed unmarked surfaces, showing any $\operatorname{Mod}(\Sigma_{g})\to\operatorname{GL}_{d}(\mathbb{C})$ with $d\leqslant 3g-3$ is forced to kill the entire Torelli subgroup.
In this article we raise Korkmaz’ lower bound to 4g34g-3 in the setting of surfaces of genus $g\geqslant 7$. We show that, when $d$ is small enough, any $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ is forced to kill the subgroup $\operatorname{SIP}_{0}(\Sigma)\leqslant\mathcal{I}(\Sigma)$ generated by the commutators $T_{a},T_{b}]$ of Dehn twists $T_{a},T_{b}\in\operatorname{PMod}(\Sigma)$ about pairs of curves $a,b\subseteq\Sigma$ intersecting at two points, with algebraic intersection pairing $\langle a,b\rangle=0$ and $\Sigma\setminus(a\cup b)$ connected.
Theorem 1 (Theorem 5.1).
Let $\Sigma$ be a surface of genus $g\geqslant 7$ and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$. If $d\leqslant 4g-4$ then $\operatorname{SIP}_{0}(\Sigma)\leqslant\ker\rho$. In particular, $d(\Sigma)\geqslant 4g-3$.
Remark.
The assumption of $g\geqslant 7$ is only used to handle the d=4g4d=4g-4 case. As a consequence, the same statement holds if we take $g\geqslant 4$ and d<4g4d<4g-4 – see Theorem 5.1. The author believes Theorem 1 should hold as stated for $g\geqslant 4$.
Unlike the subgroups $K_{\Sigma^{\prime}}^{(m)}$ from Korkmaz’ proof, the subgroup $\operatorname{SIP}_{0}(\Sigma)\leqslant\ker\rho$ from our proof remains the same regardless of $d$. It is a natural subgroup of the group $\operatorname{SIP}(\Sigma)\leqslant\mathcal{I}(\Sigma)$ generated by the so-called simple intersection maps: the commutators of twists about curves intersecting at two points and whose algebraic intersection number vanishes.
These maps were introduced by Putman in [33] as part of a generating set for the Torelli group of an unmarked genus 0 surface. Putman would then go on to use such maps in his infinite presentation of $\mathcal{I}(\Sigma_{g})$ [34]. The groups $\operatorname{SIP}(\Sigma_{g})$ and $\operatorname{SIP}(\Sigma_{g}^{1})$ were also investigated in their own right by Childers [5], who proposed a systematic study of their properties.
We also establish lower bounds for the dimensions of faithful representations of related groups. Our proofs are elementary in nature, relying mostly on well known facts about surface mapping class groups.
A new ingredient is a finer study of the commutation relations in $\operatorname{Mod}(\Sigma)$. We make use of such relations and certain families of curves to produce quotients of $F_{2}\times\cdots\times F_{2}$, the direct product of $n$ copies of a rank-$2$ free group, inside different subgroups $G\leqslant\operatorname{Mod}(\Sigma)$.
Such quotients are then used to bound the dimensions of faithful representations of certain $G\leqslant\operatorname{Mod}(\Sigma)$ by the smallest dimension of a faithful representation of $F_{2}\times\cdots\times F_{2}$. The latter was recently computed by Kionke–Schesler [21], who showed that the dimension of a faithful representation of $F_{2}\times\cdots\times F_{2}$ is $\geqslant 2n$.
Taking the following subgroups for convenience and denoting $d(G)=\min\{d|\rho:G\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})\text{ is faithful}\}$ as above, we arrive at the theorems bellow.
  1. (1)
    The Johnson subgroup $\mathcal{K}(\Sigma_{g})\leqslant\mathcal{I}(\Sigma_{g})$.
  2. (2)
    The terms $\mathcal{I}(\Sigma_{g}^{1})=\mathcal{I}^{0}(\Sigma_{g}^{1})\triangleright\mathcal{I}^{1}(\Sigma_{g}^{1})\triangleright\cdots\triangleright\mathcal{I}^{k}(\Sigma_{g}^{1})\triangleright\cdots$ of the Johnson filtration.
  3. (3)
    The pure braid group $PB_{n}$ on $n$ strands.
Theorem 2 (Corollary 3.3).
Let $g\geqslant 2$. Then $d(\mathcal{K}(\Sigma_{g}))\geqslant 2g-2$.
Theorem 3 (Corollary 3.4).
Let $g\geqslant 2$ and $k\geqslant 1$. Then $d(\mathcal{I}^{k}(\Sigma_{g}^{1}))\geqslant 2g-2$.
Theorem 4 (Corollary 3.5).
If $n$ is odd then $d(PB_{n})\geqslant n-1$. If $n$ is even then $d(PB_{n})\geqslant n-2$.
Remark.
The lower bound of n1n-1 from Theorem 4 is well known for the full braid group $B_{n}$. Dyer–Formanek–Grossman showed in [9, Proposition 2] that if $\rho:B_{n}\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is faithful then one of its irreducible subquotients must also be faithful. The irreducible representations of $B_{n}$ of dimension at most n1n-1 were classified by Formanek [11]. In particular, no irreducible representation of dimension <n1<n-1 is faithful for $n\geqslant 3$.
Using a similar strategy, we provide some evidence that greater lower bounds for the low-genus cases should lead to improvements of the lower bounds in Theorem 1.
Theorem 5 (Theorem 3.6).
Let $n\geqslant 1$ and $g\geqslant 2n$. Then $d(\Sigma_{g}^{1})\geqslant n\cdot\min\{d(E):E\text{ is a cyclic extension of }\operatorname{Mod}(\Sigma_{\lfloor\nicefrac{{g}}{{n}}\rfloor,1})\}$.
Remark.
To the best of the authors knowledge, the value of $d(\Sigma_{g,1})$ is unknown even for $g=2$. The group $\operatorname{Mod}(\Sigma_{2}^{1})$ contains a natural copy of $B_{5}$, so that $d(\Sigma_{2}^{1})\geqslant d(B_{5})$. The smallest faithful representation of $B_{5}$ known in the literature seems to be the so-called Lawrence representation $B_{5}\hookrightarrow\operatorname{GL}_{10}(\mathbb{Z}[q^{\pm 1},t^{\pm 1}])$ [1, 25]. If one assumed $d(E)\geqslant 10$ for all other cyclic extensions $E$ of $\operatorname{Mod}(\Sigma_{2,1})$ then Theorem 5 would say $d(\Sigma_{g}^{1})\geqslant 5g$, thus improving our lower bounds of 4g34g-3.
In fact, the above method is quite general, and one could imagine applying it to many other groups of interest in low-dimensional topology. Taking $G=\operatorname{PMod}(\Sigma)$ for a surface $\Sigma$ of even genus $g\geqslant 4$ even, we recover Korkmaz’ lower bound of 3g23g-2.
To go beyond this bound we instead use a different strategy. We consider a family of simple closed curves $a_{1},\ldots,a_{3g-3},b_{1},\ldots,b_{3g-3}\subseteq\Sigma$ where the curves $a_{i}$ come from a certain pants decomposition of $\Sigma_{g}$, while the curves $b_{j}$ are “complementary” to the curves $a_{i}$ – see §5 for a definition.
Given $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d$ small enough, we show that, unless $\operatorname{SIP}_{0}(\Sigma)\leqslant\ker\rho$, the matrices $M_{i}=\rho(T_{a_{i}})-1$ and $N_{j}=\rho(T_{b_{j}})-1$ satisfy the “annihilation relations”
(1.1) $N_{j}M_{i}$ $=0\iff i\neq j$ $M_{j}N_{i}$ $=0\iff i\neq j$ $M_{j}M_{i}$ $=0\;\forall i,j,$
where $T_{a}$ denotes the Dehn twist about $a\subseteq\Sigma$. This is accomplished using the disjointness relations in $\operatorname{PMod}(\Sigma)$, as well as a careful study of the eigenvalues and eigenspaces of $\rho(T_{a})$ following the work of Korkmaz, Kasahara and Kaufmann–Salter–Zhang–Zhong.
We establish lower bounds for $d$ such that we can find $M_{1},\ldots,M_{n},N_{1},\ldots,N_{n}\in M_{d}(\mathbb{C})$ satisfying (1.1). Together with the previous assertion about $M_{i}=\rho(T_{a_{i}})-1$ and $N_{j}=\rho(T_{b_{j}})-1$, such lower bounds show $\rho$ is indeed forced to kill $\operatorname{SIP}_{0}(\Sigma)$, thus concluding the proof.
1.1. Outline of the Paper
In §2 we review the theory of mapping class groups needed for the rest of the paper. This includes some cohomological calculations, used to handle the d=4g4d=4g-4 case of Theorem 1. The informed reader is invited to skip this section entirely if so inclined.
In §3 we study commutation relations in $\operatorname{PMod}(\Sigma)$. We use these relations and certain families of curves to produce quotients of $F_{2}\times\cdots\times F_{2}$ inside of different $G\leqslant\operatorname{PMod}(\Sigma)$. We then use these subgroups to establish the lower bounds from Theorem 2, Theorem 3 and Theorem 4. We then adapt this strategy to establish Theorem 5.
Still in §3, we establish a lower bound for $d$ such that we can find $M_{1},\ldots,M_{n},N_{1},\ldots,N_{n}\in M_{d}(\mathbb{C})$ satisfying (1.1) (Proposition 3.8). In §4 we study the eigenspaces of $L_{a}=\rho(T_{a})$, where $a\subseteq\Sigma$ is nonseparating and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ is low-dimensional. We establish a lower bound for the dimension of the $1$-eigenspace of $L_{a}$ (Proposition 4.2).
Finally, in §5 we conclude our proof of Theorem 1 by applying Proposition 3.8 and Proposition 4.2.
Acknowledgments
I would like to express my deep gratitude to my advisors J. Marché and M. Wolff for taking me as their student and suggesting me this problem, as well as for the incredibly helpful discussions we shared in the past year.
I should also thank F. Costantino for suggesting me to consider representations of the Torelli group. I am thankful to B. Petri for communicating me the Kaufmann–Salter–Zhang–Zhong preprint, and to G. Massuyeau for communicating me Childers’ results on the groups $\operatorname{SIP}(\Sigma_{g})$ and $\operatorname{SIP}(\Sigma_{g}^{1})$. Finally, I am grateful to J. Martel and the anonymous referees for their helpful suggestions on earlier drafts.
This work received support from the University Research School EUR-MINT (State support managed by the National Research Agency for Future Investments program bearing the reference ANR-18-EURE-0023). This work was also supported by the MathPhDInFrance International Doctoral Training Program.
2. Background Results on Mapping Class Groups
Let $\Sigma=\Sigma_{g,r}^{b}$ be the compact surface of genus $g$ with $b$ boundary components and $r$ marked points in its interior. We denote by $P$ its set of marked points. We may freely omit $b$ and $r$ from the notation when they are zero. All surfaces considered in the present paper have the form $\Sigma_{g,r}^{b}$.
The mapping class group $\operatorname{Mod}(\Sigma)=\pi_{0}\operatorname{Diff}^{+}(\Sigma,\partial\Sigma)_{P}$ is the group of orientation-preserving self-diffeomorphisms of $\Sigma$ up to isotopy, where both our diffeomorphisms and isotopies are assumed to fix the boundary point-wise and permute the marked points. The group $\operatorname{Mod}(\Sigma)$ acts on the set $P$, and the pure mapping class group $\operatorname{PMod}(\Sigma)$ is the subgroup of mapping classes acting trivially.
The braid group on $n$ strands is $B_{n}=\operatorname{Mod}(\mathbb{D}_{n})$, the mapping class group of a disk $\mathbb{D}_{n}=\Sigma_{0,n}^{1}$ with $n$ marked points. This is isomorphic to the the fundamental group of the configuration space of $n$ unordered points in a disk. The pure braid group on $n$ strands is $PB_{n}=\operatorname{PMod}(\mathbb{D}_{n})$.
In this section we collect the results from the theory of mapping class groups needed in the rest of the paper. We refer the reader to [10] for further information on mapping class groups.
Given a closed subsurface $\Sigma^{\prime}\subseteq\Sigma$ with marked points $P^{\prime}\subseteq P$, there is an induced group homomorphism $\iota:\operatorname{PMod}(\Sigma^{\prime})\to\operatorname{PMod}(\Sigma)$. Such a homomorphism needs not be injective, but we nevertheless refer to the post-composition of $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ by $\iota$ as the restriction of $\rho$ to $\operatorname{PMod}(\Sigma^{\prime})$.
2.1. Curves & Dehn Twists
Given an unoriented simple closed curve $\alpha\subseteq\Sigma$ avoiding the marked points of $\Sigma$, we denote by $a$ its free homotopy class and write $a\subseteq\Sigma$”. Here our homotopies are assumed to avoid the marked points. All curves considered in this paper are simple closed curves, unless explicitly stated otherwise.
Recall that the geometric intersection number between $a,b\subseteq\Sigma$ is the infimum
$|a\pitchfork b|=\min\{\#(\alpha\pitchfork\beta):\alpha\in a,\beta\in b\}$
of the number of times two transverse representatives of $a$ and $b$ cross each other. On the other hand, the algebraic intersection number $\langle a,b\rangle$ between $a$ and $b$ is the sum of the indices of the intersection points $x\in\alpha\pitchfork\beta$ for any $\alpha\in a$ and $\beta\in b$.
We denote by $T_{a}$ the right Dehn twist about $a$. This is the class of a diffeomorphism of $\Sigma$ supported in an annular neighborhood of $\alpha\in a$ which “winds a full turn around $\alpha$.”
It is also useful consider the twists about curves parallel to the boundary components of $\Sigma$. For example, by collapsing the boundary into a marked point we obtain a surjective group homomorphism $\operatorname{Mod}(\Sigma_{g}^{1})\to\operatorname{Mod}(\Sigma_{g,1})$. Its kernel is the subgroup generated by the Dehn twist $T_{d}$ about the boundary $d=\partial\Sigma_{g}^{1}$.
Improving results of Hatcher–Thurston [15] and Harer [14], Wajnryb [36, 2] produced a remarkable finite presentation of $\operatorname{PMod}(\Sigma)$, whose generators are given by Dehn twists and whose relations can all be explained in terms of the topology of $\Sigma$. In this paper we only need a small fragment of this result.
Theorem 2.1 (Dehn–Lickorish, [8, 26, 27]).
The group $\operatorname{PMod}(\Sigma)$ is generated by finitely many Dehn twists about nonseparating simple closed curves.
We also summarize some of the most useful relations in $\operatorname{PMod}(\Sigma)$.
  1. (1)
    The conjugation relation. Given $a\subseteq\Sigma$ and $f\in\operatorname{PMod}(\Sigma)$, $T_{f(a)}=fT_{a}f^{-1}$.
  2. (2)
    The disjointness relation. Given $a,b\subseteq\Sigma$, $T_{a}$ commutes with $T_{b}$ if and only if $|a\pitchfork b|=0$, i.e. if and only if we can find disjoint representatives for $a$ and $b$.
  3. (3)
    The braid relation. Given $a,b\subseteq\Sigma$ with $|a\pitchfork b|=1$, $T_{a}T_{b}T_{a}=T_{b}T_{a}T_{b}$.
Here a crucial observation is due: given nonseparating $a,b\subseteq\Sigma$, $T_{a}$ and $T_{b}$ are conjugate in $\operatorname{PMod}(\Sigma)$. Indeed, we can always find $f\in\operatorname{PMod}(\Sigma)$ such that $f(a)=b$, so that $fT_{a}f^{-1}=T_{b}$ by the conjugation relation. Together with Theorem 2.1, this implies the Abelianization $\operatorname{PMod}(\Sigma)^{\operatorname{ab}}=\operatorname{PMod}(\Sigma)/[\operatorname{PMod}(\Sigma),\operatorname{PMod}(\Sigma)]$ of $\operatorname{PMod}(\Sigma)$ is cyclic. In fact, $\operatorname{PMod}(\Sigma)^{\operatorname{ab}}$ vanishes when $g\geqslant 3$.
Theorem 2.2 (Powell, [32]).
Let $\Sigma$ be a surface of genus $g\geqslant 3$. Then $\operatorname{PMod}(\Sigma)$ is a perfect group.
The groups $\operatorname{PMod}(\Sigma)^{\operatorname{ab}}$ are also known in the low-genus cases. See [22, 10].
In a complementary direction, we can also consider the absence of relations between two Dehn twists.
  1. (4)
    Free subgroups. Given $a,b\subseteq\Sigma$ with $|a\pitchfork b|\geqslant 2$, $T_{a}$ and $T_{b}$ generate a rank-$2$ free group in $\operatorname{PMod}(\Sigma)$ [10, Theorem 3.14].
It turns out free subgroups are ubiquitous in $\operatorname{PMod}(\Sigma)$. This is because of the so-called Tits-alternative: a subgroup $G\leqslant\operatorname{PMod}(\Sigma)$ is either virtually Abelian or it contains free groups [30, Theorem A]. In particular, up to taking powers, the subgroup generated by two mapping classes is almost always a free group.
Theorem 2.3 (Theorem B [30]).
Let $\Sigma$ be a surface of genus $g\geqslant 2$ and $f,g\in\operatorname{PMod}(\Sigma)$. Then we can find $n,m\geqslant 1$ such that $f^{n},g^{m}$ either commute or generate a free group.
In §3 we will make use of the following result in our proof of Theorem 5.
Proposition 2.4.
Let $g\geqslant 2$ and $N,K\triangleleft\operatorname{Mod}(\Sigma_{g}^{b})$ be non-central normal subgroups. Then $N\cap K$ contains a copy of the rank-$2$ free group.
The idea is to show that any non-central $N\triangleleft\operatorname{Mod}(\Sigma_{g}^{b})$ contains pseudo-Anosov maps [^1] See [10, §13.2.3] for a definition., which is known since the mid 1980’s – see [28, Lemma 2.5]. This is also a consequence of the existence of all pseudo-Anosov normal subgroups of $\operatorname{Mod}(\Sigma_{g}^{b})$, a fact first established by Dahmani–Guirardel–Osin [7, Theorem 2.31].
2.2. The Torelli group & its Subgroups
Let $\Sigma=\Sigma_{g,r}^{b}$ be a genus $g$ surface with $r$ marked point $P\subseteq\Sigma$ and $b$ boundary components.
The natural action of $\operatorname{Diff}^{+}(\Sigma,\partial\Sigma)_{P}$ on $\Sigma$ induces a $\mathbb{Z}$-linear action of $\operatorname{Mod}(\Sigma)$ on the first homology group $H_{1}(\Sigma\setminus P;\mathbb{Z})$. The Dehn twist $T_{a}$ about $a\subseteq\Sigma$ acts by the operator
(2.1) $(T_{a})_{*}x=x+\langle a,x\rangle a\in H_{1}(\Sigma\setminus P;\mathbb{Z}),$
where $\langle\,$$\rangle$ denotes the intersection pairing. Here we view $a\subseteq\Sigma$ as an element of $H_{1}(\Sigma\setminus P;\mathbb{Z})$ by choosing an orientation of this curve. Notice, however, that (2.1) is independent of such a choice.
The Torelli subgroup of $\Sigma$, denoted $\mathcal{I}(\Sigma)\leqslant\operatorname{PMod}(\Sigma)$, is the subgroup of mapping classes acting trivially on $H_{1}(\Sigma\setminus P;\mathbb{Z})$. Some of its elements include the following maps.
  1. (1)
    Bounding pair maps. Given disjoint nonseparating curves $a,b\subseteq\Sigma$ such that $a\cup b$ bounds a closed subsurface $\Sigma^{\prime}$ with no marked points and boundary $\partial\Sigma^{\prime}=a\cup b$, the pair $(a,b)$ is called a bounding pair and $T_{a}T_{b}^{-1}$ is called a bounding pair map. By (2.1), $(T_{a})_{*}$ only depends on the homology class of $a$. Thus $T_{a}T_{b}^{-1}\in\mathcal{I}(\Sigma)$.
  2. (2)
    Separating Dehn twists. Given some separating $a\subseteq\Sigma$, $T_{a}$ is called a genus $h$ separating Dehn twists if $a$ cuts $\Sigma$ into subsurfaces of genus $h$ and $h^{\prime}$ with $h\leqslant h^{\prime}$. Since $a$ is separating, $\langle a,x\rangle=0$ for all $x\in H_{1}(\Sigma\setminus P;\mathbb{Z})$. Hence $T_{a}\in\mathcal{I}(\Sigma)$ by (2.1).
  3. (3)
    Simple intersection maps. Given $a,b\subseteq\Sigma$ with geometric intersection number $2$ and algebraic intersection number $\langle a,b\rangle=0$, the pair $(a,b)$ is called a simple intersection pair and the commutator $T_{a},T_{b}]=T_{a}T_{b}T_{a}^{-1}T_{b}^{-1}$ is called a simple intersection map. Since $\langle a,b\rangle=0$, $(T_{a})_{*}$ and $(T_{b})_{*}$ commute by (2.1), so that $T_{a},T_{b}]\in\mathcal{I}(\Sigma)$.
We can view $\Sigma$ as a subsurface of the closed unmarked genus $g$ surface $\Sigma_{g}$ by capping all boundary components with disks. Hence we can also consider the action of $\operatorname{PMod}(\Sigma)$ on $H_{\mathbb{Z}}=H_{1}(\Sigma_{g};\mathbb{Z})$, the first homology of the closed surface of the same genus. This is a rank-$2g$ free Abelian group, freely generated by the homology class of the oriented curves $a_{1},\ldots,a_{g},b_{1},\ldots,b_{g}$ in Figure 1.
$b_{1}$$a_{1}$$b_{2}$$a_{2}$$b_{g}$$a_{g}$
Figure 1. A basis for the homology of $\Sigma_{g}$.
We thus obtain a linear representation $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{2g}(\mathbb{Z})$. Its image is exactly the discrete symplectic group $\operatorname{Sp}_{2g}(\mathbb{Z})$. Indeed, the action of $\operatorname{PMod}(\Sigma)$ preserves the intersection pairing $\langle\,$$\rangle,$ which corresponds to the standard symplectic form in $H_{1}(\Sigma_{g};\mathbb{R})\cong\mathbb{R}^{2g}$.
This is called the symplectic representation $\Psi:\operatorname{PMod}(\Sigma)\to\operatorname{Sp}_{2g}(\mathbb{Z})$. More concretely, $\Psi$ is given by
(2.2) $\Psi(T_{a})x=x+\langle a,x\rangle a\in H_{\mathbb{Z}},$
not to be confused with the operators $(T_{a})_{*}$ from (2.1). When $\Sigma=\Sigma_{g}$, $\Psi(T_{a_{i}})$ and $\Psi(T_{b_{i}})$ are given by the transvections
$\Psi(T_{a_{i}})$ $=\left(\begin{array}[]{c|c|c}1&0&0\\ \hline\cr 0&\begin{matrix}1&1\\ 0&1\end{matrix}&0\\ \hline\cr 0&0&1\end{array}\right)$ $\Psi(T_{b_{i}})$ $=\left(\begin{array}[]{c|c|c}1&0&0\\ \hline\cr 0&\begin{matrix}1&0\\ -1&1\end{matrix}&0\\ \hline\cr 0&0&1\end{array}\right)$
in the basis from Figure 1, where the top-left blocks are $2(i-1)\times 2(i-1)$.
The symplectic kernel $\ker\Psi$ contains the Torelli subgroup $\mathcal{I}(\Sigma)$. This two subgroups coincide when $\Sigma$ has no marked points. Improving results of Birman [3] and Powell [32], Johnson [16] found an explicit generating set for the Torelli subgroup.
Theorem 2.5 (Johnson, [16]).
Let $\Sigma=\Sigma_{g}$ or $\Sigma=\Sigma_{g}^{1}$ be an unmarked surface of genus $g\geqslant 3$ with at most one boundary component. The Torelli subgroup $\mathcal{I}(\Sigma)$ is generated by finitely many bounding pair maps.
Now assume $\Sigma=\Sigma_{g}^{1}$ is an unmarked surface with one boundary component. By choosing a base point $\in\partial\Sigma_{g}^{1}$, the mapping class group $\operatorname{Mod}(\Sigma_{g}^{1})$ naturally acts on $\Gamma=\pi_{1}(\Sigma_{g}^{1},*)$ by group automorphisms. The Torelli subgroup can then be seen as the subgroup of mapping classes acting trivially on the Abelianization $\Gamma/[\Gamma,\Gamma]=H_{1}(\Sigma_{g}^{1};\mathbb{Z})$.
By considering the remaining terms $\Gamma_{k}=[\Gamma,\Gamma_{k-1}]$ of the lower central series of $\Gamma=\Gamma_{1}$, Johnson introduced a filtration
$\mathcal{I}(\Sigma_{g}^{1})=\mathcal{I}^{1}(\Sigma_{g}^{1})\triangleright\mathcal{I}^{2}(\Sigma_{g}^{1})\triangleright\cdots\triangleright\mathcal{I}^{k}(\Sigma_{g}^{1})\triangleright\cdots,$
known as the Johnson filtration of $\mathcal{I}(\Sigma_{g}^{1})$. Here $\mathcal{I}^{k}(\Sigma_{g}^{1})$ denotes the (normal) subgroup of mapping classes acting trivially on the characteristic quotient $\Gamma/\Gamma_{k}$.
The subgroups $\mathcal{I}^{k}(\Sigma_{g}^{1})$ are nontrivial for arbitrarily large $k$, and each $\mathcal{I}^{k}(\Sigma_{g}^{1})$ contains the $k^{\text{th}}$ term $\mathcal{I}(\Sigma_{g}^{1})_{k}$ of the lower central series of $\mathcal{I}(\Sigma_{g}^{1})=\mathcal{I}(\Sigma_{g}^{1})_{1}$. In particular, $\mathcal{I}^{2}(\Sigma_{g}^{1})\geqslant[\mathcal{I}(\Sigma_{g}^{1}),\mathcal{I}(\Sigma_{g}^{1})]$ and the quotient $\mathcal{I}(\Sigma_{g}^{1})/\mathcal{I}^{2}(\Sigma_{g}^{1})$ is Abelian. Indeed, Johnson essentially showed $\mathcal{I}(\Sigma_{g}^{1})/\mathcal{I}^{2}(\Sigma_{g}^{1})\cong{\bigwedge}^{3}H_{\mathbb{Z}}$.
The subgroup $\mathcal{K}(\Sigma_{g}^{1})=\mathcal{I}^{2}(\Sigma_{g}^{1})$ is called the Johnson subgroup of $\mathcal{I}(\Sigma_{g}^{1})$, while the projection $\tau:\mathcal{K}(\Sigma_{g}^{1})\twoheadrightarrow{\bigwedge}^{3}H_{\mathbb{Z}}$ is called the Johnson homomorphism. We may also consider $\mathcal{K}(\Sigma_{g})=\ker\tau$, where $\tau:\mathcal{I}(\Sigma_{g})\twoheadrightarrow{\bigwedge}^{3}H_{\mathbb{Z}}/H_{\mathbb{Z}}$ is given by
$\mathcal{I}(\Sigma_{g}^{1})$${\bigwedge}^{3}H_{\mathbb{Z}}$$\mathcal{I}(\Sigma_{g})$${\bigwedge}^{3}H_{\mathbb{Z}}/H_{\mathbb{Z}}.$$\scriptstyle{\tau}$$\scriptstyle{\tau}$
Here the inclusion $H_{\mathbb{Z}}\hookrightarrow{\bigwedge}^{3}H_{\mathbb{Z}}$ takes $x\in H_{\mathbb{Z}}$ to $a_{1}\wedge b_{1}\wedge x+\cdots+a_{g}\wedge b_{g}\wedge x$ for $a_{1},b_{1},\ldots,a_{g},b_{g}$ as in Figure 1. In an abuse of notation, $\mathcal{K}(\Sigma_{g})$ and $\tau:\mathcal{I}(\Sigma_{g})\twoheadrightarrow{\bigwedge}^{3}H_{\mathbb{Z}}/H_{\mathbb{Z}}$ are also called the Johnson subgroup and the Johnson homomorphism, respectively. Johnson showed that $\mathcal{K}(\Sigma_{g})$ may also be characterized as follows.
Theorem 2.6 (Johnson, [17]).
Let $g\geqslant 3$. Then $\mathcal{K}(\Sigma_{g})$ is the subgroup generated by all genus $1$ and genus $2$ separating Dehn twists.
The subgroup $\operatorname{SIP}(\Sigma)\leqslant\mathcal{I}(\Sigma)$ generated by simple intersection maps remains less well understood than $\mathcal{I}(\Sigma)$ and $\mathcal{K}(\Sigma)$. This is a normal subgroup of $\operatorname{Mod}(\Sigma)$: given a simple intersection pair $(a,b)$ and $f\in\operatorname{Mod}(\Sigma)$,
$f[T_{a},T_{b}]f^{-1}=[fT_{a}f^{-1},fT_{b}f^{-1}]=[T_{f(a)},T_{f(b)}]\in\operatorname{SIP}(\Sigma)$
by the conjugation relation.
When $\Sigma=\Sigma_{g}^{1}$, the images of simple intersection maps under the Johnson homomorphism were computed independently by Putman [35], Church [6] and Childers [5, Main Result 1]. Childers also computed the image of $\operatorname{SIP}(\Sigma_{g}^{1})$ under the so-called Birman–Craggs–Johnson homomorphism [5, Main Result 4].
We further restrict our attention to the subgroup $\operatorname{SIP}_{0}(\Sigma)\leqslant\operatorname{SIP}(\Sigma)$ generated by simple intersection maps $T_{a},T_{b}]$ where $\Sigma\setminus(a\cup b)$ is connected. This is also a normal subgroup, normally generated by $T_{a},T_{b}]$ for any simple intersection pair $(a,b)$ with $\Sigma\setminus(a\cup b)$ connected. Indeed, given any other choice of $(a^{\prime},b^{\prime})$ as above, we can find $f\in\operatorname{Mod}(\Sigma)$ with $f(a)=a^{\prime}$ and $f(b)=b^{\prime}$, implying that all generators of $\operatorname{SIP}_{0}(\Sigma)$ are conjugate in $\operatorname{Mod}(\Sigma)$.
2.3. The Work of Korkmaz
Let $\Sigma$ be a surface as above. We now describe some results due to Korkmaz which are needed in the rest of the article. We begin by his classification theorem.
The starting point of Korkmaz’ classification program is the aforementioned computation of the Abelianization of $\operatorname{PMod}(\Sigma)$. Using the braid and disjointness relations, Korkmaz [23] showed that, when $g\geqslant 2$, any linear representation $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d<2g$ factors through the Abelianization map $\operatorname{PMod}(\Sigma)\twoheadrightarrow\operatorname{PMod}(\Sigma)^{\operatorname{ab}}$.
Since $\operatorname{PMod}(\Sigma)^{\operatorname{ab}}$ vanishes for $g\geqslant 3$ (Theorem 2.2), any such representation must be trivial. Korkmaz furthermore used the same relations to show that all nontrivial $2g$-dimensional representations are conjugate to the symplectic representation $\Psi$.
Theorem 2.7 (Korkmaz, Theorems 1 & 2 [23]).
Let $\Sigma$ be a surface of genus $g\geqslant 3$ and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$. If $d\leqslant 2g$ then $\rho$ is either trivial or conjugate to $\Psi:\operatorname{PMod}(\Sigma)\to\operatorname{Sp}_{2g}(\mathbb{Z})$.
Combining this last result with Theorem 2.2, Korkmaz also established the following triviality criterion.
Lemma 2.8 (Flag triviality criterion, Lemma 7.1 [23]).
Let $\Sigma$ be a surface of genus $g\geqslant 3$ and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$. Suppose there exists a $\operatorname{PMod}(\Sigma)$-invariant flag
$0\leqslant W_{1}\leqslant W_{2}\leqslant\cdots\leqslant W_{k}=\mathbb{C}^{d}$
with $\dim W_{k}/W_{k+1}<2g$. Then $\rho$ is trivial.
As it turns out, invariant flags are pervasive. This is because of the following principle. Given $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ and $a\subseteq\Sigma$, denote $L_{a}=\rho(T_{a})$. If $a,b\subseteq\Sigma$ are disjoint then $L_{a}$ and $L_{b}$ commute, so that $L_{b}$ preserves the eigenspaces of $L_{a}$. Combining this observation with Theorem 2.1 we obtain the following.
Lemma 2.9 (Korkmaz, Lemma 4.1 [23]).
Let $\Sigma^{\prime}\subseteq\Sigma$ be closed subsurface and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$. Take $a\subseteq\Sigma\setminus\Sigma^{\prime}$. Then $E_{\lambda,k}^{a}=\ker(L_{a}-\lambda)^{k}$ is a $\operatorname{PMod}(\Sigma^{\prime})$-invariant subspace of $\mathbb{C}^{d}$. In particular, the flag
$0\leqslant E_{\lambda,1}^{a}\leqslant E_{\lambda,2}^{a}\leqslant\cdots\leqslant E_{\lambda,d}^{a}\leqslant\mathbb{C}^{d}$
is $\operatorname{PMod}(\Sigma^{\prime})$-invariant.
2.4. Twisted Cohomology
Let $\Sigma=\Sigma_{g,r}^{b}$ be the genus $g$ compact surface with $b$ boundary components and $r$ marked points. In this subsection we review some results on the cohomology of $\operatorname{PMod}(\Sigma)$. These will be used in §4 to handle the d=4g4d=4g-4 case of Theorem 1. We refer the reader to [4] for a comprehensive account of the theory of group cohomology.
Denote by $\mathbb{Z}[\operatorname{PMod}(\Sigma)]=\bigoplus_{f\in\operatorname{PMod}(\Sigma)}\mathbb{Z}f$ the group ring of $\operatorname{PMod}(\Sigma)$: the ring of (formal) integral combinations of elements in $\operatorname{PMod}(\Sigma)$, where multiplication is given by the product in $\operatorname{PMod}(\Sigma)$. Given a representation $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(R)$, we may view $R^{d}$ as a $\mathbb{Z}[\operatorname{PMod}(\Sigma)]$-module where $f\in\operatorname{PMod}(\Sigma)$ acts by $\rho(f)$. For example, the trivial $\mathbb{Z}[\operatorname{PMod}(\Sigma)]$-module is the module $\mathbb{Z}$ corresponding to the trivial homomorphism $\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{1}(\mathbb{Z})$.
Recall that, given a $\mathbb{Z}[\operatorname{PMod}(\Sigma)]$-module $M$, a map $c:\operatorname{PMod}(\Sigma)\to M$ is called an $M$-valued crossed homomorphism if $c(fg)=c(f)+f\cdot c(g)$ for all $f,g\in\operatorname{PMod}(\Sigma)$. The collection of all such maps forms an Abelian group. A crossed homomorphism $c$ is called principal if there is $m\in M$ such that $c(f)=m-f\cdot m$ for all $f$.
The first group cohomology group of $\operatorname{PMod}(\Sigma)$ with coefficients in $M$, denoted $H^{1}(\operatorname{PMod}(\Sigma);M)$, is the quotient of the group of crossed homomorphisms by the subgroup of principal crossed homomorphisms. Its elements are in one-to-one correspondence with isomorphism classes of extensions of the trivial $\mathbb{Z}[\operatorname{PMod}(\Sigma)]$-module by $M$, i.e. short exact sequences of the form
[…]
In a similar way, one may define higher cohomology groups $H^{k}(\operatorname{PMod}(\Sigma);M)$ for $k\geqslant 0$.
As a consequence of Theorem 2.2, we obtain the following computation.
Lemma 2.10.
Let $\Sigma$ be a surface of genus $g\geqslant 3$. Then $H^{1}(\operatorname{PMod}(\Sigma);\mathbb{Z})=0$.
We also consider the cohomology with coefficients in the $\mathbb{Z}[\operatorname{PMod}(\Sigma)]$-modules $H_{\mathbb{Z}}=H_{1}(\Sigma_{g};\mathbb{Z})$ and $H_{\mathbb{C}}=H_{1}(\Sigma_{g};\mathbb{C})$ corresponding to the symplectic representation $\Psi:\operatorname{PMod}(\Sigma)\to\operatorname{Sp}_{2g}(\mathbb{Z})$. We pay special attention to the surfaces $\Sigma_{g}^{1}$ and $\Sigma_{g,1}$. Such groups were first computed by Morita [31].
Theorem 2.11 (Morita, Proposition 6.4 [31]).
If $g\geqslant 2$ then $H^{1}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}})\cong H^{1}(\operatorname{Mod}(\Sigma_{g,1});H_{\mathbb{Z}})\cong\mathbb{Z}$.
A standard argument with the universal coefficient theorem shows $H^{1}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{C}})\cong\mathbb{C}$ for $g\geqslant 2$. This implies that all nontrivial extensions of the $H_{\mathbb{C}}$ by $\mathbb{C}$ are isomorphic as $\mathbb{Z}[\operatorname{Mod}(\Sigma_{g}^{1})]$-modules. Kasahara [18] showed that, up to dualizing, any nontrivial $(2g+1)$-dimensional $\mathbb{Z}[\operatorname{PMod}(\Sigma_{g,r}^{b})]$-module is isomorphic to one such extension. Here we make use of a homological lemma of his.
Lemma 2.12 (Kasahara, Theorem 4.2 [18]).
Let $c:\operatorname{Mod}(\Sigma_{g}^{1})\to H_{\mathbb{C}}$ be a crossed homomorphism. Given $a\subseteq\Sigma_{g}^{1}$ nonseparating, $c(T_{a})=\lambda\cdot a$ for some $\lambda\in\mathbb{C}$.
Morita’s result was later generalized by Kawazumi [20], who computed the (stable) higher cohomology groups of $\operatorname{Mod}(\Sigma_{g}^{1})$ with coefficients in tensor powers of $H_{\mathbb{Z}}$ in terms of the so called twisted Miller–Mumford–Morita classes $\hat{\kappa}_{P}\in H^{*}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}}^{\otimes n})$ associated to weighted partitions $P$ of the set $\{1,\ldots,n\}$. See [20] for definitions.
Theorem 2.13 (Kawazumi, Theorem 1.B [20]).
For $k\leqslant\nicefrac{{g}}{{2}}-n$,
$H^{k}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}}^{\otimes n})=\bigoplus_{\begin{subarray}{c}P\in\mathcal{P}_{n}\\ \ell+\deg\hat{\kappa}_{P}=k\end{subarray}}H^{\ell}(\operatorname{Mod}(\Sigma_{g}^{1});\mathbb{Z})\smile\hat{\kappa}_{P}.$
Here $\mathcal{P}_{n}$ denotes the set of weighted partitions of $\{1,\ldots,n\}$ and
$H^{\ell}(\operatorname{Mod}(\Sigma_{g}^{1});\mathbb{Z})\smile\hat{\kappa}_{P}=\{\xi\smile\hat{\kappa}_{P}:\xi\in H^{\ell}(\operatorname{Mod}(\Sigma_{g}^{1});\mathbb{Z})\},$
where $\xi\smile\hat{\kappa}_{P}\in H^{k}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}}^{\otimes n})$ is the cup product of $\xi$ with $\hat{\kappa}_{P}$.
When $n=2$, the twisted Miller–Mumford–Morita classes take two forms: the classes $\alpha_{i}\in H^{2i}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}}^{\otimes 2})$ where $i\geqslant 0$, and the classes $\beta_{ij}\in H^{2i+2j-2}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}}^{\otimes 2})$ where $i,j\geqslant 1$, corresponding to partitions of $\{1,2\}$ into one and two subsets, respectively. We thus obtain the following computation.
Corollary 2.14.
For $g\geqslant 6$, $H^{1}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}}^{\otimes 2})=0$.
Proof.
Since $1\leqslant\nicefrac{{g}}{{2}}-2$, Theorem 2.13 says $H^{1}(\operatorname{Mod}(\Sigma_{g}^{1});H_{\mathbb{Z}}^{\otimes 2})=H^{1}(\operatorname{Mod}(\Sigma_{g}^{1});\mathbb{Z})\smile\alpha_{0}$. Now recall from Lemma 2.10 that $H^{1}(\operatorname{Mod}(\Sigma_{g}^{1});\mathbb{Z})=0$ already for $g\geqslant 3$. ∎
3. Commutation Relations in the Mapping Class Group
Recall that the commuting graph of a group $G$ is the graph $\Gamma(G)$ whose vertices are elements of $G$, where $g,h\in G$ are joined by an edge if and only if they commute. First defined by Harvey, the curve graph of a surface $\Sigma$ is the graph $\mathcal{C}(\Sigma)$ whose vertices are homotopy classes of essential simple closed curves in $\Sigma$, where $a,b\subseteq\Sigma$ are joined by an edge if and only if we can find disjoint representatives.
The latter is a Gromov-hyperbolic graph on which $\operatorname{PMod}(\Sigma)$ acts by isometries [29]. Given the disjointness relations from §2.1, $\mathcal{C}(\Sigma)$ and $\Gamma(\operatorname{PMod}(\Sigma))$ are related by means of the embedding $\mathcal{C}(\Sigma)\hookrightarrow\Gamma(\operatorname{PMod}(\Sigma))$ taking $a\subseteq\Sigma$ to its Dehn twist $T_{a}$.
The starting point of the present article was to consider the disjointness relations in $\operatorname{PMod}(\Sigma)$ induced by a family curves $a_{1},\ldots,a_{n},b_{1},\ldots,b_{n}\subseteq\Sigma$ satisfying
(3.1) $|a_{i}\pitchfork a_{j}|$ $=0\;\;\forall i\neq j$ $|b_{i}\pitchfork b_{j}|$ $=0\;\;\forall i\neq j$ $|a_{i}\pitchfork b_{j}|$ $=0\;\;\forall i\neq j$ $|a_{i}\pitchfork b_{i}|$ $\geqslant 1\;\;\forall i,$
where $|a\pitchfork b|$ denotes the geometric intersection number between $a$ and $b$.
These correspond to copies of the graph $\Delta_{n}$ from Figure 2 inside $\mathcal{C}(\Sigma)$. In terms of $\Gamma(\operatorname{PMod}(\Sigma))$, such a family of curves translates to the relations
(3.2) $T_{a_{i}}T_{a_{j}}$ $=T_{a_{j}}T_{a_{i}}\text{ for all }i,j$ $T_{b_{i}}T_{b_{j}}$ $=T_{b_{j}}T_{b_{i}}\text{ for all }i,j$ $T_{a_{i}}T_{b_{j}}$ $=T_{b_{j}}T_{a_{i}}\iff i\neq j.$
These relations can be further refined as follows.
$a_{1}$$b_{1}$$a_{2}$$b_{2}$$a_{n-1}$$b_{n-1}$$a_{n}$$b_{n}$
Figure 2. The graph $\Delta_{n}$: the vertices $a_{1},\ldots,a_{n}$ and $b_{1},\ldots,b_{n}$ form two disjoint $n$-cliques, while the vertices $a_{i}$ and $b_{j}$ are connected by an edge if and only if $i\neq j$.
Assume $a_{1},\ldots,a_{n},b_{1},\ldots,b_{n}\subseteq\Sigma$ is a family as above, with the first $k$ pairs $(a_{i},b_{i})$ intersecting exactly once and the remaining $\ell=n-k$ pairs intersecting at $\geqslant 2$ points. By the braid relation, $T_{a_{i}}$ and $T_{b_{i}}$ generate a copy of the braid group $B_{3}$ inside $\operatorname{PMod}(\Sigma)$ for $i\leqslant k$. Likewise, $T_{a_{i}}$ and $T_{b_{i}}$ generate a rank-$2$ free group $F_{2}$ for $i\geqslant k+1$.
Combined with (3.2), these relations then imply there is a a well defined homomorphism from
$G_{k\ell}=\underbrace{B_{3}\times\cdots\times B_{3}}_{k\text{ times}}\times\underbrace{F_{2}\times\cdots\times F_{2}}_{\ell\text{ times}}$
onto the subgroup generated by $T_{a_{1}},\ldots,T_{a_{n}},T_{b_{1}},\ldots,T_{b_{n}}$. What is more, the projection of each direct factor of $G_{k\ell}$ onto its image in $\operatorname{PMod}(\Sigma)$ is an isomorphism.
More generally, one can produce quotients of $G_{k\ell}$ inside a subgroup $G\leqslant\operatorname{PMod}(\Sigma)$ by considering the subgroups of $G$ consisting of mapping classes supported on disjoint closed subsurfaces $S_{1},\ldots,S_{n}\subseteq\Sigma$, each one corresponding to one of the factors of $G_{k\ell}$. It is thus natural to expect the dimension of a faithful representation $G\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ to be related to the minimal dimension of a faithful representation of $G_{k\ell}$.
The latter was recently estimated by Kionke–Schesler [21].
Proposition 3.1 (Kionke–Schesler, Theorem 3 [21]).
Suppose $H_{1},\ldots,H_{n}$ are non-solvable groups and $\rho:H_{1}\times\cdots\times H_{n}\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is a faithful linear representation. Then $d\geqslant 2n$.
Their proof is short and elementary, making clever use of well known facts about the representation theory of direct products. Here we make use of a slightly more general version of their statement, although our proof is really an adaptation of their argument.
Proposition 3.2.
Suppose $H_{1},\ldots,H_{n}$ are non-solvable groups and $\pi:H_{1}\times\cdots\times H_{n}\twoheadrightarrow H$ is a surjective group homomorphism such that $\pi\!\restriction_{H_{i}}:H_{i}\to H$ is injective for all $i$. Let $\rho:H\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ be a faithful linear representation. Then $d\geqslant 2n$.
Proof.
Consider $\rho\circ\pi:H_{1}\times\cdots\times H_{n}\to\operatorname{GL}_{d}(\mathbb{C})$ and take a maximal $(H_{1}\times\cdots\times H_{n})$-invariant flag
$0=W_{0}\leqslant W_{1}\leqslant\cdots\leqslant W_{p+1}=\mathbb{C}^{d},$
so that the action $\rho_{i}:H_{1}\times\cdots\times H_{n}\to\operatorname{GL}(W_{i+1}/W_{i})$ of $H_{1}\times\cdots\times H_{n}$ on each successive quotient is irreducible. In a basis adapted to this flag,
(3.3) $\rho(\pi(h))=\begin{pmatrix}\rho_{1}(h)&*&\cdots&*\\ 0&\rho_{2}(h)&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\rho_{p}(h)\end{pmatrix}$
for all $h\in H_{1}\times\cdots\times H_{n}$.
Now each $W_{i+1}/W_{i}$ may be decomposed as a tensor product $W_{i+1}/W_{i}=W_{i,1}\otimes\cdots\otimes W_{i,n}$, where $\rho_{ij}:H_{j}\to\operatorname{GL}(W_{ij})$ is an irreducible representation and $\rho_{i}(h_{1},\ldots,h_{n})=\rho_{i,1}(h_{1})\otimes\cdots\otimes\rho_{i,n}(h_{n})$ – see, for example, [24, Proposition 2.3.23]. In this setting,
$\dim W_{i+1}/W_{i}=(\dim W_{i,1})\cdots(\dim W_{i,n})\geqslant 2^{\#\{j:\dim W_{ij}\geqslant 2\}}\geqslant 2\cdot\#\{j:\dim W_{ij}\geqslant 2\},$
for $2^{k}\geqslant 2k$ for all integers $k\geqslant 0$. In particular, $d=\dim W_{1}/W_{0}+\cdots+\dim W_{p+1}/W_{p}\geqslant 2\cdot\#\{(i,j):\dim W_{ij}\geqslant 2\}$.
If $\dim W_{ij}=1$ then $\rho_{ij}(h_{j})$ is a scalar operator for all $h_{j}\in H_{j}$, so that $\rho_{i}(1,\ldots,h_{j},\ldots,1)=1\otimes\cdots\otimes\rho_{ij}(h_{j})\otimes\cdots\otimes 1$ is also a scalar operator. Assume we can find $j\leqslant n$ with $\dim W_{ij}=1$ for all $i$. Then the matrix $\rho(\pi(1,\ldots,h_{j},\ldots,1))$ is upper triangular with respect to the basis from (3.3) for all $h_{j}\in H_{j}$. Since the group of upper triangular matrices is solvable, it follows $\pi(H_{i}^{(k)})\leqslant\ker\rho$ for some $k$.
But $H_{i}$ is non-solvable and so $\pi(H_{i}^{(k)})\cong H_{i}^{(k)}\neq 1$, contradicting the assumption that $\rho$ is faithful. This means that, for each $j\leqslant n$ we can find $i$ such that $\dim W_{ij}\geqslant 2$. Hence $d\geqslant 2\cdot\#\{(i,j):\dim W_{ij}\geqslant 2\}\geqslant 2n$, as desired. ∎
Our proofs of Theorem 2, Theorem 3 and Theorem 4 are direct applications of Proposition 3.2.
Corollary 3.3 (Theorem 2).
Let $g\geqslant 2$ and suppose $\rho:\mathcal{K}(\Sigma_{g})\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is a faithful representation of the Johnson kernel. Then $d\geqslant 2g-2$.
Proof.
Consider the curves $a_{1},\ldots,a_{g-1},c_{1},\ldots,c_{g-1}\subseteq\Sigma_{g}$ as in Figure 3(a) and take $b_{i}=T_{c_{i}}(a_{i})$. For each $i$, the curves $a_{i}$ and $c_{i}$ intersect twice. The geometric intersection number of $a_{i}$ and $b_{i}$ is thus $4=2^{2}$. It is also clear all other pairs of curves in the above family are disjoint.
$c_{1}$$a_{1}$$c_{g-1}$$a_{g-1}$
(a) $\Sigma=\Sigma_{g}$
$c_{1}$$a_{1}$$c_{g-1}$$a_{g-1}$
(b) $\Sigma=\Sigma_{g}^{1}$
Figure 3. The curves $a_{1},\ldots,a_{g-1},c_{1},\ldots,c_{g-1}\subseteq\Sigma$.
The curves $a_{i}$ and $b_{j}$ are all separating. The discussion above then implies that the subgroup of $\mathcal{K}(\Sigma_{g})$ generated by $T_{a_{1}},\ldots,T_{a_{g-1}},T_{b_{1}},\ldots,T_{b_{g-1}}$ is a quotient of $G_{0,g-1}$, the direct product of g1g-1 copies of a rank-$2$ free group $F_{2}$. What is more, the projection of each $F_{2}$-factor of $G_{0,g-1}$ onto its image in $\mathcal{K}(\Sigma_{g})$ is an isomorphism. The result thus follows from Proposition 3.2. ∎
To prove Theorem 3, we pass to the derived subgroups $\mathcal{I}(\Sigma_{g}^{1})^{(k)}=[\mathcal{I}(\Sigma_{g}^{1})^{(k-1)},\mathcal{I}(\Sigma_{g}^{1})^{(k-1)}]$ of $\mathcal{I}(\Sigma_{g}^{1})=\mathcal{I}(\Sigma_{g}^{1})^{(1)}$.
Corollary 3.4 (Theorem 3).
Let $\Sigma_{g}^{1}$ be the unmarked genus $g$ surface with one boundary component. Suppose $g\geqslant 2$. If $\rho:\mathcal{I}(\Sigma_{g}^{1})^{(k)}\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is faithful then $d\geqslant 2g-2$. In particular, if $\rho:\mathcal{I}^{k}(\Sigma_{g}^{1})\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is faithful then $d\geqslant 2g-2$.
Proof.
For the first claim, consider the curves $a_{1},\ldots,a_{g-1},c_{1},\ldots,c_{g-1}\subseteq\Sigma_{g}^{1}$ from Figure 3(b) and take $b_{i}=T_{c_{i}}(a_{i})$. As in the proof of Corollary 3.3, the subgroup of $\mathcal{I}(\Sigma_{g}^{1})$ generated by $T_{a_{1}},\ldots,T_{a_{g-1}},T_{b_{1}},\cdots,T_{b_{g-1}}$ is a quotient of $G_{0,g-1}=F_{2}\times\cdots\times F_{2}$.
In particular, $\mathcal{I}(\Sigma_{g}^{1})^{(k)}$ contains a quotient of the $k^{\text{th}}$ derived subgroup $G_{0,g-1}^{(k)}=F_{2}^{(k)}\times\cdots\times F_{2}^{(k)}$. What is more, the projection of each $F_{2}^{(k)}$-factor onto its image in $\mathcal{I}(\Sigma_{g}^{1})^{(k)}$ is an isomorphism. Since $F_{2}$ is non-solvable, so is $F_{2}^{(k)}$. The result thus follows from Proposition 3.2.
For the second claim, it suffices to observe $\mathcal{I}(\Sigma_{g}^{1})^{(k)}\leqslant\mathcal{I}(\Sigma_{g}^{1})_{k}\leqslant\mathcal{I}^{k}(\Sigma_{g}^{1})$, where $\mathcal{I}(\Sigma_{g}^{1})_{k}=[\mathcal{I}(\Sigma_{g}^{1}),\mathcal{I}(\Sigma_{g}^{1})_{k-1}]$ are the terms of the lower central series of $\mathcal{I}(\Sigma_{g}^{1})=\mathcal{I}(\Sigma_{g}^{1})_{1}$. ∎
Corollary 3.5 (Theorem 4).
Suppose $\rho:PB_{n}\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is faithful. If $n$ is odd then $d\geqslant n-1$. If $n$ is even then $d\geqslant n-2$.
Proof.
Recall $PB_{n}=\operatorname{PMod}(\mathbb{D}_{n})$ is the pure mapping class group of a disk with $n$ marked points. Given $m<n$, the natural map $PB_{m}\to PB_{n}$ is injective, so that we may regard $PB_{m}$ as subgroup of $PB_{n}$.
In particular, if $n$ is even, we can pass to the subgroup $PB_{n-1}\leqslant PB_{n}$. We may thus assume $n=2k+1$ for some $k\geqslant 0$. Furthermore, the result is clearly true for $n\leqslant 3$. We can thus assume $k\geqslant 2$.
$b_{1}$$b_{2}$$b_{k}$$a_{1}$$a_{2}$$a_{k}$
Figure 4. The curves $a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\subseteq\mathbb{D}_{n}$.
In that case, consider $a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\subseteq\mathbb{D}_{n}$ as in Figure 4. It follows from the discussion above that the subgroup generated by $T_{a_{1}},\ldots,T_{a_{k}},T_{b_{1}},\ldots,T_{b_{k}}\in\operatorname{PMod}(\mathbb{D}_{n})$ is a quotient of $G_{0,k}$, the direct product of $k$ copies of $F_{2}$. What is more, the projection of each $F_{2}$-factor onto its image in $\operatorname{PMod}(\Sigma_{0,n}^{1})$ is an isomorphism. Proposition 3.2 then says $d\geqslant 2k=n-1$, as desired. ∎
Recall $d(G)$ denotes the smallest $d$ such that one can find a faithful $G\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$, and $d(\Sigma)=d(\operatorname{PMod}(\Sigma))$. Let $\Sigma_{g,1}$ be the closed genus $g$ surface with a single marked point. By replacing the number $2$ by $\min\{d(E):E\text{ is a cyclic extension of }\operatorname{Mod}(\Sigma_{\lfloor\nicefrac{{g}}{{n}}\rfloor,1})\}$ in the proof of Proposition 3.2 we obtain Theorem 5.
Theorem 3.6 (Theorem 5).
Let $n\geqslant 1$ and $g\geqslant 2n$. Then $d(\Sigma_{g}^{1})\geqslant n\cdot\min\{d(E):E\text{ is a cyclic extension of }\operatorname{Mod}(\Sigma_{\lfloor\nicefrac{{g}}{{n}}\rfloor,1})\}$.
Proof.
Let $\rho:\operatorname{Mod}(\Sigma_{g}^{1})\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ be a faithful representation and $g^{\prime}=\lfloor\nicefrac{{g}}{{n}}\rfloor$. Take $d_{\operatorname{min}}=\min\{d(E):E\text{ is a cyclic extension of }\operatorname{Mod}(\Sigma_{g^{\prime},1})\}$. We want to show $d\geqslant n\cdot d_{\operatorname{min}}$. By passing to a smaller subsurface $\Sigma_{n\cdot g^{\prime}}^{1}\subseteq\Sigma_{g}^{1}$ if necessary, we may assume $g=n\cdot g^{\prime}$ with $g^{\prime}\geqslant 2$. In that case, we may view $\Sigma_{g}^{1}$ as an $(n+1)$-hole sphere with $n$ copies $S_{1},\ldots,S_{n}\subseteq\Sigma_{g}^{1}$ of $\Sigma_{g^{\prime}}^{1}$ attached along their boundaries.
The natural maps $\operatorname{Mod}(S_{i})\to\operatorname{Mod}(\Sigma_{g}^{1})$ are injective, so that we may regard $\operatorname{Mod}(S_{i})$ as a subgroup of $\operatorname{Mod}(\Sigma_{g}^{1})$. Since $S_{i}\cap S_{j}=\emptyset$ for $i\neq j$, there is a well-defined homomorphism $\pi:\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n})\to\operatorname{Mod}(\Sigma_{g}^{1})$ with $\pi\!\restriction_{\operatorname{Mod}(S_{j})}:\operatorname{Mod}(S_{j})\to\operatorname{Mod}(\Sigma_{g}^{1})$ injective for all $i$.
Take a maximal $(\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n}))$-invariant flag
(3.4) $0=W_{0}\leqslant W_{1}\leqslant\cdots\leqslant W_{p+1}=\mathbb{C}^{d},$
so that the action $\rho_{i}:\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n})\to\operatorname{GL}(W_{i+1}/W_{i})$ of $\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n})$ on each successive quotient is irreducible. Set $W_{i+1}/W_{i}=W_{i,1}\otimes\cdots\otimes W_{i,n}$, where $\rho_{ij}:\operatorname{Mod}(S_{j})\to\operatorname{GL}(W_{ij})$ is an irreducible representation and $\rho_{i}(f_{1},\ldots,f_{n})=\rho_{i,1}(f_{1})\otimes\cdots\otimes\rho_{i,n}(f_{n})$ [24, Proposition 2.3.23].
For each $j$, we may regard $\rho_{ij}$ as a representation of $\operatorname{Mod}(\Sigma_{g^{\prime}}^{1})\cong\operatorname{Mod}(S_{j})$. Let us show that, for each $j\leqslant n$, we can find $i$ with $\dim W_{ij}\geqslant d_{\operatorname{min}}$. In that case,
$\begin{split}d&=\dim W_{1}/W_{0}+\cdots+\dim W_{p+1}/W_{p}\\ &\geqslant d_{\operatorname{min}}^{\#\{j:\dim W_{1,j}\geqslant d_{\operatorname{min}}\}}+\cdots+d_{\operatorname{min}}^{\#\{j:\dim W_{p,j}\geqslant d_{\operatorname{min}}\}}\\ &\geqslant d_{\operatorname{min}}\cdot\#\{(i,j):\dim W_{ij}\geqslant d_{\operatorname{min}}\}\\ &\geqslant n\cdot d_{\operatorname{min}},\end{split}$
as desired.
Fix $j\leqslant n$. Assume at first we can find $i$ such that $\ker\rho_{ij}\leqslant\operatorname{Mod}(\Sigma_{g^{\prime}}^{1})$ is central. Since $g^{\prime}\geqslant 2$, the center of $\operatorname{Mod}(\Sigma_{g^{\prime}}^{1})$ is generated by $T_{d}$, the Dehn twist about the boundary curve $d=\partial\Sigma_{g^{\prime}}^{1}$. As in §2.1, the quotient $\operatorname{Mod}(\Sigma_{g^{\prime}}^{1})/T_{d}\cong\operatorname{Mod}(\Sigma_{g^{\prime},1})$ is the mapping class group of the closed genus $g^{\prime}$ surface with one marked point. In particular, $\operatorname{Mod}(\Sigma_{g^{\prime}}^{1})$ is a central extension of $\operatorname{Mod}(\Sigma_{g^{\prime},1})$ by $\langle T_{d}\rangle\cong\mathbb{Z}$.
If $\rho_{ij}$ is faithful then $\dim W_{ij}\geqslant d(\Sigma_{g^{\prime}}^{1})\geqslant d_{\operatorname{min}}$ by definition. We may thus assume $\ker\rho_{ij}\neq 1$, in which case it is freely generated by a power $T_{d}^{k_{i}}$ of $T_{d}$ with $k_{i}\geqslant 1$. In that case, $\operatorname{Mod}(\Sigma_{g^{\prime}}^{1})/\ker\rho_{ij}=\operatorname{Mod}(\Sigma_{g^{\prime}}^{1})/T_{d}^{k_{i}}$ is a central extension of $\operatorname{Mod}(\Sigma_{g^{\prime},1})$ by $\mathbb{Z}/k_{i}$. In particular, $\dim W_{ij}\geqslant d_{\operatorname{min}}$ once again.
We are left to consider the case where, for some $j$, $\ker\rho_{ij}$ is not central for all $i$. Let us show this situation cannot happen. Denoting $K=\ker\rho_{1,j}\cap\cdots\cap\ker\rho_{p,j}$, it is clear $K$ acts on $\mathbb{C}^{d}$ by operators which are upper triangular with respect to a basis adapted to the flag from (3.4). Hence its $k^{\text{th}}$ derived subgroup $K^{(k)}$ lies in $\ker\rho$ for large enough $k$. But $K$ contains a free subgroup by Proposition 2.4. This implies $K^{(k)}\neq 1$, which contradicts the assumption $\rho$ is faithful. We are done. ∎
We now focus our attention on the faithful representations of $\operatorname{PMod}(\Sigma)$. A simple count shows that the maximal size of a family $a_{1},\ldots,a_{n},b_{1},\ldots,b_{n}\subseteq\Sigma_{g}$ as in (3.1) whose pairwise geometric intersection numbers are $\leqslant 2$ is 2n=3g22n=3g-2 or 2n=3g32n=3g-3, depending on whether $g$ is even or odd, respectively. As a consequence, Proposition 3.2 thus recovers lower bounds similar to Korkmaz’.
These families can be obtained by viewing $\Sigma_{g}$ as a $g$-holed sphere attached to $1$-holed tori $H_{1},\ldots,H_{g}\subseteq\Sigma_{g}$. The $g$ first pairs $(a_{i},b_{i})$ are taken as $a_{i},b_{i}\subseteq H_{i}$ intersecting once. The remaining pairs can be obtained by subdividing the $g$-holed sphere into g2g-2 pairs of pants and combining them into $4$-holed spheres $S_{1},\ldots,S_{\lfloor\nicefrac{{g-2}}{{2}}\rfloor}\subseteq\Sigma_{g}$. We then choose $a_{g+i},b_{g+i}\subseteq S_{i}$ intersecting twice.
To move beyond Korkmaz’ bound of 3g23g-2 we instead use a different strategy. We consider a family of curves $a_{1},\ldots,a_{3g-3},b_{1},\ldots,b_{3g-3}\subseteq\Sigma$ where the curves $a_{i}$ come from a certain pants decomposition of $\Sigma_{g}$, while the curves $b_{j}$ are, in some sense, “complementary” to the curves $a_{i}$. Unlike the curves in (3.1), the curve $b_{j}$ is allowed to intersect $b_{i}$ twice for $i\neq j$. See §5 for a proper definition.
Take $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d$ small enough. In §4 and §5 we will show that, unless $\rho$ kills the generators of $\operatorname{SIP}_{0}(\Sigma)$, the matrices $M_{i}=\rho(T_{a_{i}})-1$ and $N_{j}=\rho(T_{b_{j}})-1$ satisfy the “annihilation relations”
(3.5) $N_{j}M_{i}$ $=0\iff i\neq j$ $M_{j}N_{i}$ $=0\iff i\neq j$ $M_{j}M_{i}$ $=0\;\forall i,j.$
We emphasize the relations $N_{j}N_{i}=0$ are not part of (3.5).
We now establish a simple lower bound for $d$ such that we can find matrices $M_{1},\ldots,M_{n},N_{1},\ldots,N_{n}\in M_{d}(\mathbb{C})$ satisfying (3.5). This will be used in §5 to show that $\rho$ is, in fact, forced to kill $\operatorname{SIP}_{0}(\Sigma)$.
Lemma 3.7.
Let $M_{1},\ldots,M_{n},N_{1},\ldots,N_{n}\in M_{d}(\mathbb{C})$ be operators such that $N_{j}M_{i}=0$ if an only if $i\neq j$. Then $\dim\sum_{i}\operatorname{range}M_{i}\geqslant n$.
Proof.
We proceed by induction in $n$. The base case $n=1$ is clear. Now suppose the theorem holds for a given $n$ and let us show the same holds for $n+1$. Given $M_{1},\ldots,M_{n+1},N_{1},\ldots,N_{n+1}$ as above, it follows from the induction hypothesis that
$\dim\sum_{i\leqslant n}\operatorname{range}M_{i}\geqslant n.$
We now claim one can find $v\in\operatorname{range}M_{n+1}$ with $v\notin\sum_{i\leqslant n}\operatorname{range}M_{i}$, so that $\dim\sum_{i\leqslant n+1}\operatorname{range}M_{i}\geqslant 1+\dim\sum_{i\leqslant n}\operatorname{range}M_{i}\geqslant n+1$.
Indeed, since $N_{n+1}M_{n+1}\neq 0$, there is $w\in\mathbb{C}^{n}$ with $N_{n+1}M_{n+1}w\neq 0$. On the other hand, $N_{n+1}(M_{1}w_{1}+\cdots+M_{n}w_{n})=N_{n+1}M_{1}w_{1}+\cdots N_{n+1}M_{n}w_{n}=0$ for all $w_{1},\ldots,w_{n}\in\mathbb{C}^{n}$. In other words, $N_{n+1}\!\restriction_{\sum_{i\leqslant n}\operatorname{range}M_{i}}=0$ and thus $v=M_{n+1}w\notin\sum_{i\leqslant n}\operatorname{range}M_{i}$. This concludes the inductive step. ∎
Proposition 3.8.
Let $M_{1},\ldots,M_{n},N_{1},\ldots,N_{n}\in M_{d}(\mathbb{C})$ be nonzero operators subject to the annihilation relations (3.5). Then $\dim\sum_{i}(\operatorname{range}M_{i}+\operatorname{range}N_{i})\geqslant 3n-d$. In particular, $2d\geqslant 3n$.
Proof.
It is clear from Lemma 3.7 that $\dim\sum_{i}\operatorname{range}M_{i}\geqslant n$ and $\dim\sum_{i}\operatorname{range}N_{i}\geqslant n$. Let us show that $\dim\left(\sum_{i}\operatorname{range}M_{i}\right)\cap\left(\sum_{i}\operatorname{range}N_{i}\right)\leqslant d-n$, so that
$\begin{split}\dim\sum_{i}(\operatorname{range}M_{i}+\operatorname{range}N_{i})&=\dim\sum_{i}\operatorname{range}M_{i}+\dim\sum_{i}\operatorname{range}N_{i}\\ &-\dim\left(\sum_{i}\operatorname{range}M_{i}\right)\cap\left(\sum_{i}\operatorname{range}N_{i}\right)\\ &\geqslant 2n-(d-n)\\ &=3n-d.\end{split}$
Since $M_{i}M_{j}=0$ for all $i$ and $j$, $\left(\sum_{i}\operatorname{range}M_{i}\right)\cap\left(\sum_{i}\operatorname{range}N_{i}\right)\leqslant\bigcap_{i}\ker M_{i}=\ker\Phi$, where $\Phi=\bigoplus_{i}M_{i}:\mathbb{C}^{d}\to\bigoplus_{i}\operatorname{range}M_{i}$. By the second relation in (3.5), we can find $w_{i}\in\mathbb{C}^{n}$ such that $M_{i}N_{i}w_{i}\neq 0$. On the other hand, $M_{j}N_{i}=0$ for $j\neq i$ and, in particular, $M_{j}N_{i}w_{i}=0$. Hence $\Phi(v_{i})\neq 0$ lies in the copy of $\operatorname{range}M_{i}$ inside of the codomain of $\Phi$ for $v_{i}=N_{i}w_{i}$.
Choosing one such $w_{i}$ for each $i=1,\ldots,n$ we get that the vectors $\Phi(v_{1}),\ldots,\Phi(v_{n})$ are linearly independent, so that $\operatorname{rank}\Phi\geqslant n$. Hence $\dim\ker\Phi\leqslant d-n$, as desired. ∎
4. Eigenspaces of $T_{a}$
Fix some $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d\leqslant 4g-4$. In this section we study the $1$-eigenspace of $L_{a}=\rho(T_{a})$ for some nonseparating $a\subseteq\Sigma$. We establish a lower bound for the dimension of the $1$-eigenspace of $L_{a}$.
As a consequence, we obtain the fact the matrices $M_{i}=L_{a_{i}}-1$ and $N_{i}=L_{b_{i}}-1$ associated to the aforementioned family $a_{1},\ldots,a_{3g-3},\ldots,b_{1},\ldots,b_{3g-3}\subseteq\Sigma$ satisfy the first two relations in (3.5) – see Corollary 4.3. This will be used in §5 to apply Proposition 3.8 to the matrices $M_{i}$ and $N_{j}$ as above.
Given $a\subseteq\Sigma$, we denote the $\lambda$-eigenspace of $L_{a}$ by $E_{\lambda}^{a}$. We also take $E_{\lambda,k}=\ker(L_{a}-\lambda)^{k}$, so that $E_{\lambda,1}^{a}=E_{\lambda}^{a}$ and $E_{\lambda,d}^{d}$ is the generalized $\lambda$-eigenspace of $L_{a}$. Recall from §2.1 that the Dehn twists about nonseparating $a,b\in\Sigma$ are conjugate in $\operatorname{PMod}(\Sigma)$. In particular, $L_{a}\sim L_{b}$. We may thus pass from one nonseparating curve to the next when performing our analysis.
We will call $\rho$ unipotent if $1$ is the only eigenvalue of $L_{a}$ for some (and hence all) nonseparating $a\subseteq\Sigma$. Establishing the unipotency of low-dimensional representations is a crucial step in the classification theorems of Korkmaz, Kasahara and Kaufmann–Salter–Zhang–Zhong. This is summarized in the following proposition.
Proposition 4.1 (Kaufmann–Salter–Zhang–Zhong, Proposition 6.1 [19]).
Let $\Sigma$ be a surface of genus $g\geqslant 4$ and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d\leqslant 4g-3$. Given $a\subseteq\Sigma$ nonseparating, the only eigenvalue of $L_{a}$ is $1$.
Building on the work of Kaufmann–Salter–Zhang–Zhong, we establish a lower bound for the following dimension of the $1$-eigenspace of $L_{a}$.
Proposition 4.2.
Let $\Sigma$ be a genus $g$ surface and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ be nontrivial. Suppose either of the following conditions are met:
  1. (1)
    $g\geqslant 4$ and d<4g4d<4g-4, or
  2. (2)
    $g\geqslant 7$ and $d\leqslant 4g-4$.
If $a\subseteq\Sigma$ is nonseparating then $\dim E^{a}_{1}>2g-2$.
Corollary 4.3.
Let $\Sigma$ be a genus $g$ surface and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ satisfying either (1) or (2) from Proposition 4.2. Take $a,b\subseteq\Sigma$ disjoint with $a$ nonseparating. Then $(L_{a}-1)(L_{b}-1)=0$.
Proof of Corollary 4.3.
Denote by $\Sigma_{a}$ the surface obtained from $\Sigma$ by cutting across $a$.
The result clearly holds for trivial $\rho$. We may thus assume $\rho$ is nontrivial. We can find a basis for $\mathbb{C}^{d}$ under which
$\rho(f)=\left(\begin{array}[]{c|c}\rho_{1}(f)&*\\ \hline\cr 0&\bar{\rho}(f)\end{array}\right)$
for all $f\in\operatorname{PMod}(\Sigma_{a})$, where the top-left and bottom-right blocks correspond to the action of $f$ on $E^{a}_{1}$ and $\mathbb{C}^{d}/E^{a}_{1}$, respectively.
Now since $\rho$ is nontrivial, $\dim E_{1}^{a}>2g-2$ by Proposition 4.2 and thus $\dim\mathbb{C}^{d}/E^{a}_{1}<2g-2$. It follows from Theorem 2.7 that $\bar{\rho}(f)=1$. Given any $b\subseteq\Sigma$ disjoint from $a$, $b\subseteq\Sigma_{a}$ and so we may write
$L_{b}-1=\left(\begin{array}[]{c|c}L_{b}\!\restriction_{E^{b}_{1}}-1&*\\ \hline\cr 0&0\end{array}\right)$
in this basis. In particular, $\operatorname{range}(L_{b}-1)\leqslant E^{a}_{1}=\ker(L_{a}-1)$. ∎
We now review some results needed for the proof of Proposition 4.2.
Lemma 4.4 (Jordan inequalities).
Let $A\in M_{d}(\mathbb{C})$ and $\lambda\in\mathbb{C}$. Consider the flag
$0=E_{\lambda,0}\leqslant E_{\lambda,1}\leqslant E_{\lambda,2}\leqslant\cdots\leqslant E_{\lambda,d},$
where $E_{\lambda,k}=\ker(A-\lambda)^{k}$. Then $\dim E_{\lambda,k+1}/E_{\lambda,k}\leqslant\dim E_{\lambda,k}/E_{\lambda,k-1}$ for all $k=0,1,\ldots,d-1$.
Proof.
The map $A-\lambda$ induces an injective linear map $(A-\lambda):E_{\lambda,k+1}/E_{\lambda,k}\hookrightarrow E_{\lambda,k}/E_{\lambda,k-1}$ for all $k$. ∎
Lemma 4.5 (Korkmaz, Lemma 4.3 [23]).
Let $\Sigma$ be a surface of genus $g\geqslant 2$ and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$. Fix two nonseparating curves $a,b\subseteq\Sigma$ intersecting at a single point and suppose $E_{\lambda}^{a}=E_{\lambda}^{b}$. Then $E_{\lambda}^{a}$ is a $\operatorname{PMod}(\Sigma)$-invariant subspace.
We are now ready to prove Proposition 4.2.
Proof of Proposition 4.2.
Take some nontrivial $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d\leqslant 4g-4$ as above, where $\Sigma$ is a surface of genus $g\geqslant 4$. Let $a\subseteq\Sigma$ be nonseparating. We know from Proposition 4.1 that $1$ is the only eigenvalue of $L_{a}$. Let $\Sigma^{\prime}\cong\Sigma_{g-1}^{1}$ be a subsurface as in Figure 5.
$\Sigma^{\prime}$bbaa
Figure 5. The subsurface $\Sigma^{\prime}\cong\Sigma_{g-1}^{1}$.
Suppose by contradiction $\dim E_{1}^{a}\leqslant 2g-2$. First, assume $\dim E_{1}^{a}<2g-2$ and consider the $\operatorname{Mod}(\Sigma^{\prime})$-invariant flag
$0\leqslant E_{1}^{a}=E_{1,1}^{a}\leqslant E_{1,2}^{a}\leqslant\cdots\leqslant E_{1,d}^{a}=\mathbb{C}^{d},$
where $E_{1,k}^{a}=\ker(L_{a}-1)^{k}$ as above. By the Jordan inequalities (Lemma 4.4), $\dim E_{1,k+1}^{a}/E_{1,k}^{a}\leqslant\dim E_{1}^{a}<2g-2$ for all $k$. It thus follows from the flag triviality criterion (Lemma 2.8) that the restriction of $\rho$ to $\operatorname{Mod}(\Sigma^{\prime})$ is trivial. But then $L_{a}\sim L_{c}=1$ for any nonseparating $c\subseteq\Sigma^{\prime}$, contradicting the assumption $\dim E_{1}^{a}<2g-2$.
It remains to show $\dim E_{1}^{a}\neq 2g-2$. Assume by contradiction $\dim E_{1}^{a}=2g-2$ and denote by $\rho_{1}:\operatorname{Mod}(\Sigma^{\prime})\to\operatorname{GL}(E_{1}^{a})$ and $\bar{\rho}:\operatorname{Mod}(\Sigma^{\prime})\to\operatorname{GL}(\mathbb{C}^{d}/E_{1}^{a})$ the actions of $\operatorname{Mod}(\Sigma^{\prime})$ on $E_{1}^{a}$ and $\mathbb{C}^{d}/E_{1}^{a}$, respectively. It follows from Theorem 2.7 that $\rho_{1}$ and $\bar{\rho}$ are either trivial or conjugate to the symplectic representation $\Psi:\operatorname{Mod}(\Sigma^{\prime})\to\operatorname{Sp}_{2g-2}(\mathbb{Z})\leqslant\operatorname{GL}(H_{\mathbb{C}})$, where $H_{\mathbb{C}}=H_{1}(\Sigma_{g-1};\mathbb{C})$.
We consider three separate cases.
Case 1. Assume $\rho_{1}$ is trivial. In this case, $L_{c}\!\restriction_{E_{1}^{a}}=\rho_{1}(T_{c})=1$ and thus $E_{1}^{a}\leqslant E_{1}^{c}$ for all nonseparating $c\subseteq\Sigma^{\prime}$. Since $L_{a}$ and $L_{c}$ are conjugate, this implies $E_{1}^{a}=E_{1}^{c}$. By the same token, $E_{1}^{b}=E_{1}^{c}=E_{1}^{a}$ for $b$ as in Figure 5. Now Lemma 4.5 implies $E_{1}^{a}=E_{1}^{b}$ is $\operatorname{PMod}(\Sigma)$-invariant.
We abuse the notation and denote by $\rho_{1}:\operatorname{PMod}(\Sigma)\to\operatorname{GL}(E_{1}^{a})$ and $\bar{\rho}:\operatorname{PMod}(\Sigma)\to\operatorname{GL}(\mathbb{C}^{d}/E_{a}^{1})$ the actions of $\operatorname{PMod}(\Sigma)$ on $E_{1}^{a}$ and $\mathbb{C}^{d}/E_{a}^{1}$, respectively. In that case, $\rho_{1}$ and $\bar{\rho}$ are both trivial by Theorem 2.7. The flag triviality criterion (Lemma 2.8) applied to the flag $0\leqslant E_{1}^{a}\leqslant\mathbb{C}^{d}$ thus implies $\rho$ is trivial, a contradiction.
Case 2. Assume $\rho_{1}\sim\Psi$ and $\bar{\rho}$ is trivial. In this case, we can find a basis for $\mathbb{C}^{d}$ under which
$\rho(f)=\left(\begin{array}[]{c|c}\Psi(f)&\begin{matrix}c_{1}(f)&c_{2}(f)&\cdots&c_{d-2g+2}(f)\end{matrix}\\ \hline\cr 0&1\end{array}\right)$
for all $f\in\operatorname{Mod}(\Sigma^{\prime})$. It is not hard to check the maps $c_{k}:\operatorname{Mod}(\Sigma^{\prime})\to E_{1}^{a}\cong H_{\mathbb{C}}$ are crossed homomorphisms.
Now Lemma 2.12 implies that, given $c\subseteq\Sigma^{\prime}$ nonseparating, $c_{k}(c)=\mu_{k}\cdot c$ for some $\mu_{k}\in\mathbb{C}$. By tweaking the above basis, we can find a second basis for $\mathbb{C}^{d}$ under which
$L_{c}=\left(\begin{array}[]{c|c}\Psi(T_{c})&\begin{matrix}\mu\cdot c&0&\cdots&0\end{matrix}\\ \hline\cr 0&1\end{array}\right)$
for some $\mu\in\mathbb{C}$. Hence $\operatorname{codim}E_{1}^{a}=\operatorname{codim}E_{1}^{c}\leqslant 2$ by (2.2), a contradiction for $d>2g$.
We may thus assume $d\leqslant 2g$, in which case Theorem 2.7 says $\rho$ is either trivial or conjugate to the symplectic representation $\Psi:\operatorname{Mod}(\Sigma)\to\operatorname{Sp}_{2g}(\mathbb{Z})$. The former contradicts the hypothesis $\rho$ is nontrivial, so $\rho\sim\Psi$. But then $\dim E_{1}^{a}=2g-1$ by (2.2), contradicting the assumption $\dim E_{1}^{a}=2g-2$.
Case 3. Finally, assume $\rho_{1}\sim\bar{\rho}\sim\Psi$. This last case is only possible if d=4g4d=4g-4, which is only relevant to our proof when $g\geqslant 7$. We thus assume d=4g4d=4g-4 and $g\geqslant 7$ from now on.
We regard $\mathbb{C}^{4g-4}$ as a $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$-module, where $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$ denotes the group ring of $\operatorname{Mod}(\Sigma^{\prime})$ and $f\in\operatorname{Mod}(\Sigma^{\prime})$ acts on $\mathbb{C}^{d}$ by $\rho(f)$, as in §2.4. In this case, $\mathbb{C}^{4g-4}$ is an extension of $H_{\mathbb{C}}=H_{1}(\Sigma_{g-1};\mathbb{C})$ by $H_{\mathbb{C}}$. This means $\mathbb{C}^{4g-4}$ fits into a short exact sequence of the form
[…]
Such extensions are classified by the group $\operatorname{Ext}^{1}_{\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]}(H_{\mathbb{C}},H_{\mathbb{C}})=\operatorname{Ext}^{1}_{\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]}(H_{\mathbb{Z}},H_{\mathbb{Z}})\otimes_{\mathbb{Z}}\mathbb{C}$, where $H_{\mathbb{Z}}=H_{1}(\Sigma_{g-1};\mathbb{Z})$.
On the one hand, $\operatorname{Ext}^{1}_{\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]}(H_{\mathbb{Z}},H_{\mathbb{Z}})=H^{1}(\operatorname{Mod}(\Sigma^{\prime});\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}}))$ [4, Proposition 2.2]. Here $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$ acts on $\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})$ by $f\cdot A=\Psi(f)\circ A\circ\Psi(f)^{-1}$ for all $f\in\operatorname{Mod}(\Sigma^{\prime})$ and $A\in\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})$. Hence $\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})\cong H_{\mathbb{Z}}^{*}\otimes_{\mathbb{Z}}H_{\mathbb{Z}}$ as $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$-modules, where $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$ acts on $H_{\mathbb{Z}}^{*}=\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},\mathbb{Z})$ via $f\cdot A=A\circ\Psi(f)^{-1}$.
On the other hand, the intersection pairing $\langle\,,\rangle:H_{\mathbb{Z}}\times H_{\mathbb{Z}}\to\mathbb{Z}$ induces a $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$-module isomorphism $H_{\mathbb{Z}}^{*}\cong H_{\mathbb{Z}}$, so that $H_{\mathbb{Z}}^{*}\otimes_{\mathbb{Z}}H_{\mathbb{Z}}\cong H_{\mathbb{Z}}^{\otimes 2}$. Since $g-1\geqslant 6$, $\operatorname{Ext}^{1}_{\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]}(H_{\mathbb{Z}},H_{\mathbb{Z}})\cong H^{1}(\operatorname{Mod}(\Sigma^{\prime});H_{\mathbb{Z}}^{\otimes 2})=0$ by Corollary 2.14. This implies the sequence (4.1) splits.
We can thus find a basis for $\mathbb{C}^{4g-4}$ under which
$\rho(f)=\left(\begin{array}[]{c|c}\Psi(f)&0\\ \hline\cr 0&\Psi(f)\end{array}\right)$
for all $f\in\operatorname{Mod}(\Sigma^{\prime})$. Taking $f=T_{c}$ for some nonseparating $c\subseteq\Sigma^{\prime}$, we can see $\dim E_{1}^{a}=\dim E_{1}^{c}=4g-6$, a contradiction. ∎
5. Lower Bounds for Faithful Representations of the Mapping Class Group
In this section we conclude our proof of Theorem 1. Let $\Sigma$ be a genus $g$ surface, possible with boundary components and marked points. We embed $\Sigma$ in $\Sigma_{g}$ by capping the boundary components with disks.
Given $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d\leqslant 4g-4$, our goal is showing $\operatorname{SIP}_{0}(\Sigma)\leqslant\ker\rho$. As mentioned before, our strategy is to apply Proposition 3.8 to the family of matrices $M_{i}=\rho(T_{a_{i}})-1$ and $N_{i}=\rho(T_{b_{i}})-1$ associated with $\rho$, where $a_{1},\ldots,a_{3g-3},b_{1},\ldots,b_{3g-3}\subseteq\Sigma$ are obtained from a certain pants decomposition of $\Sigma_{g}$. We begin by defining the curves $a_{i},b_{j}\subseteq\Sigma$.
Consider a trivalent graph $\Gamma_{g}$ with 2g22g-2 vertices given as follows. We start by arranging 2g22g-2 vertices uniformly in a circle, so that, for each vertex we draw, we also draw its antipode. We then join each vertex in the circle with the two adjacent vertices and its antipode, as in Figure 6. We embed $\Gamma_{g}$ in $3$-space as to avoid edge intersections.
Figure 6. The graph $\Gamma_{g}$ for $g=4$, with the neighborhood $U_{i}$ of $e_{i}$ highlighted in green.
The graph $\Gamma_{g}$ is connected and has 3g33g-3 edges $e_{1},e_{2},\ldots,e_{3g-3}$. What is more, for each such edge $e_{i}$ we can find a small “double-Y-shaped” neighborhood $U_{i}\subseteq\Gamma_{g}$ of $e_{i}$ such that $\Gamma_{g}\setminus U_{i}$ is still connected, as in Figure 7. By thickening $\Gamma_{g}$ we obtain a genus $g$ handlebody with boundary $\Sigma_{g}$.
For each $e_{i}$, let $a_{i}\subseteq\Sigma_{g}$ be a meridian around $e_{i}$. These curves form a pants decomposition of $\Sigma_{g}$. By thickening the neighborhood $U_{i}$ of $e_{i}$ we obtain a neighborhood $S_{i}$ of $a_{i}$ which is a $4$-holed sphere. This neighborhood may also be obtained by gluing two adjacent pairs of pants along their common boundary component $a_{i}$. Let $b_{i}\subseteq\Sigma_{g}$ be as in Figure 7.
$S_{i}$$e_{i}$$U_{i}$
(a) The neighborhoods $U_{i}$ and $S_{i}$.
$b_{i}$$a_{i}$
(b) $a_{i},b_{i}\subseteq S_{i}$
Figure 7. Definition of the curves $a_{i}$ and $b_{i}$.
We can find representatives of $a_{i}$ and $b_{i}$ lying on $\Sigma\subseteq\Sigma_{g}$ and avoiding its marked points, which we denote by $a_{i},b_{i}\subseteq\Sigma$ as well. Although such curves depend on a choice of representative of $a_{i},b_{i}\subseteq\Sigma_{g}$, this choice is inconsequential to us. Since $\Gamma_{g}\setminus U_{i}$ is connected, so is $\Sigma_{g}\setminus S_{i}$. In particular, $a_{i}$ and $b_{i}$ are nonseparating.
It is clear the curves $a_{i}$ and $a_{j}$ are disjoint for $i\neq j$. While $a_{i}$ is disjoint from $b_{j}$ for all $i\neq j$, a careful reader might notice the curves $b_{i}$ and $b_{j}$ can intersect each other nontrivially when $e_{i}$ and $e_{j}$ are adjacent edges of $\Gamma_{g}$. We will see this will not pose a problem to our proof.
We are now ready to prove Theorem 1. Recall from §2.2 that the subgroup $\operatorname{SIP}_{0}(\Sigma)$ generated by the simple intersection maps $T_{a},T_{b}]$ with $\Sigma\setminus(a\cup b)$ connected is a normal subgroup of $\mathcal{I}(\Sigma)$, normally generated by any such generator.
Theorem 5.1 (Theorem 1).
Let $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$. If $g\geqslant 4$ and d<4g4d<4g-4 then $\operatorname{SIP}_{0}(\Sigma_{g})\leqslant\ker\rho$. Moreover, if $g\geqslant 7$ and d=4g4d=4g-4 then $\operatorname{SIP}_{0}(\Sigma_{g})\leqslant\ker\rho$.
Proof.
Given $a\subseteq\Sigma_{g}$, denote $L_{a}=\rho(T_{a})$ and $E_{1}^{a}=\ker(L_{a}-1)$ as above. Suppose by contradiction $\ker\rho$ does not contain $\operatorname{SIP}_{0}(\Sigma)$. This means $L_{a}$ and $L_{b}$ do not commute for some (and hence all) nonseparating $a,b\subseteq\Sigma$ intersecting twice with $\Sigma\setminus(a\cup b)$ is connected.
Take $a_{1},\ldots,a_{3g-3},b_{1},\ldots,b_{3g-g}\subseteq\Sigma$ as above. We claim that the matrices
(5.1) $M_{i}$ $=L_{a_{i}}-1$ $N_{j}$ $=L_{b_{j}}-1$
satisfy the conditions in (3.5). In that case, it follows from Proposition 3.8 that $2d\geqslant 9g-9$, a contradiction for $d\leqslant 4g-4$. To establish the claim, notice $\Gamma_{g}\setminus U_{i}$ is connected. This implies $\Sigma_{g}\setminus S_{i}$ is connected. Hence so is $\Sigma\setminus(a_{i}\cup b_{i})$. In particular, $\operatorname{SIP}_{0}(\Sigma)$ is normally generated by $T_{a_{i}},T_{b_{i}}]$ for any $i$.
It is clear from Corollary 4.3 that $N_{j}M_{i}=M_{j}N_{i}=0$ for $i\neq j$ and $M_{j}M_{i}=0$ for all $i,j$. Since the curves $b_{i}$ and $b_{j}$ can intersect each other nontrivially, we cannot conclude $N_{i}N_{j}=0$. This fact is, however, immaterial to our proof, as the relations $N_{i}N_{j}=0$ are not part of the hypothesis of Proposition 3.8.
On the other hand, by assumption, $L_{a_{i}}$ and $L_{b_{i}}$ do not commute. Hence $M_{i}$ and $N_{i}$ don’t commute. In particular, $M_{i}N_{i}\neq 0$ or $N_{i}M_{i}\neq 0$. But $M_{i}N_{i}=0\iff N_{i}M_{i}=0$. Indeed, the pairs $(T_{a_{i}},T_{b_{i}})$ and $(T_{b_{i}},T_{a_{i}})$ are conjugate in $\operatorname{PMod}(\Sigma)$: we can find $f\in\operatorname{PMod}(\Sigma)$ with $f(a_{i})=b_{i}$ and $f(b_{i})=a_{i}$, so that $fT_{a_{i}}f^{-1}=T_{b_{i}}$ and $fT_{b_{i}}f^{-1}=T_{a_{i}}$. The operators $M_{i}N_{i}$ and $N_{i}M_{i}$ are thus conjugated by $\rho(f)\in\operatorname{GL}_{d}(\mathbb{C})$. We are done. ∎
Disclaimer
Co-funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.